# RSHazeDiff: A Unified Fourier-aware Diffusion Model for Remote Sensing Image Dehazing

### v1

@@ Abstract
Haze severely degrades the visual quality of remote sensing images and hampers the performance of automotive navigation, intelligent monitoring, and urban management. The emerging denoising diffusion probabilistic model (DDPM) exhibits the significant potential for dense haze removal with its strong generation ability. Since remote sensing images contain extensive small-scale texture structures, it is important to effectively restore image details from hazy images. However, current wisdom of DDPM fails to preserve image details and color fidelity well, limiting its dehazing capacity for remote sensing images. In this paper, we propose a novel unified Fourier-aware diffusion model for remote sensing image dehazing, termed RSHazeDiff. From a new perspective, RSHazeDiff explores the conditional DDPM to improve image quality in dense hazy scenarios, and it makes three key contributions. First, RSHazeDiff refines the training phase of diffusion process by performing noise estimation and reconstruction constraints in a coarse-to-fine fashion. Thus, it remedies the unpleasing results caused by the simple noise estimation constraint in DDPM. Second, by taking the frequency information as important prior knowledge during iterative sampling steps, RSHazeDiff can preserve more texture details and color fidelity in dehazed images. Third, we design a global compensated learning module to utilize the Fourier transform to capture the global dependency features of input images, which can effectively mitigate the effects of boundary artifacts when processing fixed-size patches. Experiments on both synthetic and real-world benchmarks validate the favorable performance of RSHazeDiff over multiple state-of-the-art methods. Source code will be released at https://github.com/jm-xiong/RSHazeDiff
@@ I Introduction
REMOTE sensing (RS) images captured under hazy environments suffer from significant quality degradation, resulting in the performance deterioration of subsequent high-level tasks (e.g., image segmentation [1], road recognition [2]). Therefore, image dehazing is a necessary preprocessing step to promote the robustness of RS applications.
As image dehazing is a typical ill-posed problem, early works develop various prior assumptions to generate haze-free images [3, 4, 5, 6, 7, 8, 9]. These hand-crafted prior assumptions aim to regularize the solution space of linear inverse problems. The well-known prior knowledge contains the dark channel prior (DCP) [3] and color attenuation prior (CAP) [4]. These methods effectively improve the overall visual quality of degraded images. However, predefined prior assumptions are only applied in specific scenarios, leading to severe color shift and undesired artifacts in practice.
@@
Recently, various learning-based dehazing approaches have been proposed, and they may outperform prior-based methods. Specifically, early efforts [10, 11, 12] exploit convolutional neural networks (CNNs) to estimate the transmission maps or the airlight, and then generate sharp images via atmospheric scattering model (ASM) [13]. However, the cumulative error and unreliable prior assumptions often cause color distortions or haze residuals. Since then, considerable end-to-end dehazing approaches [14, 15, 16, 17, 18, 19] focus on directly generating haze-free images from the corresponding hazy images. With the advantage of generative models, most dehazing methods based on generative adversarial networks (GANs) [20, 21, 22, 23, 24, 25] regularize data distribution of dehazed images for haze removal. However, the training of GANs is unstable and susceptible to mode corruption, so GAN-based algorithms fail to achieve promising dehazing results.
@@
Diffusion model is an emerging generative paradigm and successful in various low-level vision tasks [26, 27, 28, 29]. With the superior modeling data distribution ability, diffusion models have achieved impressive performance in image generation. Moreover, the diffusion model is not subjected to unstable training and mode-collapse as GAN-based methods. Motivated by the strong generation ability of diffusion models, we leverage the conditional DDPM [30] for remote sensing image dehazing (RSID). However, there still exist several problems that limit the dehazing performance of DDPM. First, DDPM only uses simple noise-estimation constraints to optimize the network while ignoring image-level refinement. That simple constraining noise may introduce chaotic content and undesired impurity into restored images. Second, DDPM cannot utilize the potential structure and statistical information in the forward iterative process, limiting its better sampling capacity. In early sampling steps, the images captured in thick hazy scenarios make it difficult to complement the structure and color features, whereas the forward iterative process preserves relatively well-established information. By exploiting the beneficial information in the forward iterative, DDPM can better recover image details and color information (see Fig. 1(d)). Third, although the patch-based DDPM enables size-agnostic processing and sampling acceleration, it yields evident block artifacts in that the patch-based strategy overlooks the global semantics of all patches. As shown in Fig. 1(e)), introducing the global compensation learning module is beneficial to eliminate boundary artifacts and produce more visually pleasing images.
@@
To overcome the aforementioned limitations, we propose a novel DDPM-based network for RSID, termed RSHazeDiff. Specifically, we regard the conditional DDPM as the basic architecture. Similar to the existing conditional strategy, RSHazeDiff adopts the degraded images as conditions to provide effective distribution guidance in the sampling process, preventing content diversity with the stochastic sampling process. In addition, we develop a phased training strategy (PTS) to effectively remedy the unsatisfactory results caused by simple noise-estimation constraint in the training phase. Concretely, we first train the diffusion model with the noise-estimation constraint, and then further optimize it with the reconstruction constraint. The reconstruction constraint aims to constrain the sampled restored results with its corresponding ground-truth images. Furthermore, to take advantage of latent structure and statistical information in the forward iterative process, we design a Fourier-aware iterative refinement (FIR) module. In each sampling step, FIR excavates the style/semantic information from the amplitude/phase components in Fourier domain for finer sampling results. Referring to the patch-based image processing in [31], we leverage the patch-based image processing strategy to ensure unlimited-size image dehazing. Nevertheless, separately dealing with each patch may ignore the global dependency features of all patches, leading to evident artifacts in the final result. In view of this, we utilize a global compensated learning module (GCL) to capture global dependency features and then incorporate GCL and local reverse denoising process into a unified format to produce high-quality output. Comprehensive experiments verify the superiority of RSHazeDiff on both synthetic and real-world RS hazy images, especially in dense hazy scenes.
@@
  • The global compensated learning (GCL) module is introduced to eliminate evident block artifacts resulting from the patch-based image processing scheme. GCL can capture the global dependency features in RS images to assure the semantic consistency of all patches.
@@ II Related Work
This section briefly reviews RSID algorithms and the diffusion model for image restoration. Existing RSID algorithms are classified into two categories: prior-based methods and learning-based methods.
@@ II-A Remote Sensing Image Dehazing
Prior-based methods. These methods [5, 6, 7, 8, 9] employ the hand-crafted priors to restore the dehazed images. For instance, Li et al. [5] develop a homomorphic filtering and sphere model improved DCP for nonhomogeneous haze and thin cloud removal. Xu et al. [6] design an iterative dehazing method named IDeRS, which introduces the concept of “virtual depth” for evaluating Earth surface coverings. Shen et al. [7] propose a spatial-spectral adaptive method for RSID, which corrects the ambient light and transmission estimation from the spatial and spectral dimensions. Han et al. [9] can effectively remove the haze of RS hazy images by exploiting multiscale guided filtering to decompose the input images into base layers and detail layers for final output. Although these prior-based methods achieve impressive results, they rely on prior assumptions in specific scenes, making the dehazing results easily violated when the assumptions are not satisfied.
Learning-based methods. With the success of deep learning, researchers have focused on learning-based RSID algorithms [18, 32, 33, 25, 19, 34, 35, 36]. Similar to natural image dehazing, some RSID methods explore deep neural networks to estimate haze parameters in ASM. For example, Chi et al. [35] develop a trinity model called Trinity-Net to incorporate prior knowledge into the gradient-guided Swin Transformer [37] for haze parameters estimation. The other works directly convert hazy images to haze-free images in an end-to-end manner. Among them, Li and Chen [18] design a coarse-to-fine dehazing network called FCTF-Net, in which the first stage exploits an encoder–decoder architecture to obtain coarse dehazing results and then refines them in second stage. Zhang and Wang [19] propose a dynamic collaborative inference learning framework (DCIL) to handle haze removal of dense RS hazy images. Zheng et al. [25] exploit an enhanced attention-guide GAN-based network named Dehaze-AGGAN to restore haze-free remote sensing images in an unsupervised manner. However, few works focus on applying the diffusion model to image dehazing task, whereas the advanced generation ability of diffusion model makes it possible to better recover haze-free images.
@@ II-B Diffusion Models for Image Restoration
Diffusion-based generative models [38, 30, 39] progressively corrupt images by adding noise, and iteratively denoise from a noise distribution to a sharp image. With the superior generative capability, diffusion-based models have achieved remarkable performance in image restoration, such as image super-resolution [26], inpainting [27], shadow removal [28] and deblurring [29]. Saharia et al. [26] utilize conditional diffusion model for image super-resolution and achieve more competitive results than GAN-based methods. With similar condition strategy, Saharia et al. [27] propose a unified framework to handle image-to-image translation tasks by incorporating distorted images as the condition input. Rombach et al. [40] apply diffusion models in latent space to improve sampling efficiency. Considering that current diffusion model architectures mainly focus on fixed-sized image restoration, Ozan et al. [31] propose a patch-based diffusion model for unlimited-size image restoration. However, it fails to capture global information when operating locally on fixed-sized patches, which yields apparent block artifacts. To address this issue, we design a global compensated learning module to capture long-range context in frequency domain.
@@ II-C Fourier Transform in Deep Learning
Recently, the frequency information extracted through Fourier transform has attracted increasing attention in deep learning. A line of works utilizes it to improve representation ability of neural networks [41, 42, 43, 44]. Among them, Xu et al. [41] demonstrate that the learning in frequency domain facilitates neural networks more efficiently and effectively. Wang et al. [43] reveal the relationship between the generalization of CNNs and the high-frequency components of images. Besides, Zhou et al. [44] propose deep Fourier up-sampling for multi-scale modeling, which achieves impressive performance on multiple computer vision tasks in a plug-and-play manner. The other line of work applies the Fourier transform on many computer vision tasks [45, 46, 47, 48]. For instance, Yang et al. [45] develop an effective phase consistency constraint to align source-target domains flexibly for domain adaptation semantic segmentation. Fuoli et al. [47] adopt Fourier domain losses to better recover the high-frequency content in image super-resolution. Yu et al. [48] explore the haze degradation property from a frequency domain perspective, and have achieved success in image dehazing by this property. However, no works attempt to exploit the potential of the Fourier spectral components in diffusion models. To our knowledge, we are the first to take the Fourier spectral components as important prior knowledge during iterative sampling steps of diffusion model, better preserving image details and colors.
@@ IV-A Overview
Fig. 2 illustrates the overall architecture of RSHazeDiff, which consists of three main components: a local reverse denoising process, a global compensated learning module, and a feature fusion module. The local reverse denoising process leverages the Fourier-aware condition diffusion model to perform local patch-level restoration, and an input hazy image is randomly cropped with several patches as its conditions. The detailed training and sampling process of Fourier-aware conditional diffusion model is exhibited in Fig. 3, and its pseudocodes are depicted in Alg. 1 and Alg. 2. Note that we crop the whole image into overlapping patches as the input of Fourier-aware conditional diffusion model by using the grid-like arranged parsing scheme [31]. In turn, we merge all patches with the dictionary of their locations and adopt the mean estimated noise for overlapping pixels [31] (lines 8,9,11 in Alg. 2). GCL module adopts the Fast Fourier Block (FFB) as the basic unit of U-net architecture to capture global feature representations. The structure of FFB is shown Fig. 4. Finally, the feature fusion module fuses the global-local feature representations to produce the final dehazing results.
@@ IV-B Fourier-aware Iterative Refinement
DDPM-based image restoration approaches tend to tackle the forward diffusion process and reverse denoising process in sequence. The reverse process progressively sample with variable $\epsilon_{0}\left(x_{t},\bar{x},t\right)$ predicted in the forward process. Apart from that, the two processes are independent to a certain extent. Nevertheless, each iteration in the forward process can provide beneficial guided features to the sampling process. As depicted in [50], the amplitude and phase components in Fourier space correspond to low-level statistics and high-level semantic information of input images. This characteristic can be extended in image dehazing task. The amplitude spectrum contains more hazy degradation properties (e.g., illumination and contrast), while the phase spectrum preserves more texture structure of images. In early sampling steps, dense hazy images suffer from severe structure and color information loss, whereas the forward iterative process preserves pretty well-established information. Inspired by that, we design a Fourier-aware iterative refinement module, which excavates the structure and color guided representation from the forward process in Fourier domain to refine each transition of sampling process. The well-designed FIR module can help diffusion model to generate more natural and realistic images.
The detailed structure of FIR is illustrated in Fig. 3. Note that the $x_{0}^{(i)}$ represents $i$-th patch cropped from the clear image $x_{0}$. we first transition step from a patch $x_{t-1}^{(i)}$ to $x_{t}^{(i)}$ through Eq. 1, and then sample from $\hat{x}_{t}^{(i)}$ to $\hat{x}_{t-1}^{(i)}$ through Eq. 4. Then, images $x_{t-1}^{(i)}$ and $\hat{x}_{t-1}^{(i)}$ are decomposed to amplitude and phase components in Fourier domain by Fast Fourier Transform (FFT) [51]. Next, we use the phase spectrum of $x_{t-1}^{(i)}$ and integrate the partial amplitude spectrum of $x_{t-1}^{(i)}$ into amplitude component of $\hat{x}_{t-1}^{(i)}$. Finally, we apply the inverse Fast Fourier Transform (IFFT) to transform the final fused results $\hat{x}_{t-1}^{(i)}$ back to spatial domain image. Given a single channel image $x\in\mathcal{R}^{H\times W}$, its Fourier transform $\mathcal{F}$ as a complex component can be expressed as:

### v2

@@ Abstract
Haze severely degrades the visual quality of remote sensing images and hampers the performance of road extraction, vehicle detection, and traffic flow monitoring. The emerging denoising diffusion probabilistic model (DDPM) exhibits the significant potential for dense haze removal with its strong generation ability. Since remote sensing images contain extensive small-scale texture structures, it is important to effectively restore image details from hazy images. However, current wisdom of DDPM fails to preserve image details and color fidelity well, limiting its dehazing capacity for remote sensing images. In this paper, we propose a novel unified Fourier-aware diffusion model for remote sensing image dehazing, termed RSHazeDiff. From a new perspective, RSHazeDiff explores the conditional DDPM to improve image quality in dense hazy scenarios, and it makes three key contributions. First, RSHazeDiff refines the training phase of diffusion process by performing noise estimation and reconstruction constraints in a coarse-to-fine fashion. Thus, it remedies the unpleasing results caused by the simple noise estimation constraint in DDPM. Second, by taking the frequency information as important prior knowledge during iterative sampling steps, RSHazeDiff can preserve more texture details and color fidelity in dehazed images. Third, we design a global compensated learning module to utilize the Fourier transform to capture the global dependency features of input images, which can effectively mitigate the effects of boundary artifacts when processing fixed-size patches. Experiments on both synthetic and real-world benchmarks validate the favorable performance of RSHazeDiff over state-of-the-art methods. Source code will be released at https://github.com/jm-xiong/RSHazeDiff
@@ I Introduction
REMOTE sensing (RS) records the surface information from high altitudes in a non-contact manner, providing a broader synoptic view for traffic monitoring with zero impact on traffic flow. Unfortunately, RS images captured by optical remote sensors inevitably suffer from noticeable degradation under hazy conditions. Such hazy images hamper visual tasks to extract traffic-related information from images with the help of intelligent transportation systems, such as road extraction [1, 2], vehicle detection [3, 4], traffic flow monitoring [5, 6], and autonomous driving [7]. RS image dehazing improves the overall visibility of degraded images and restores important traffic information, which reduces the risk of image misinterpretation. RS image dehazing is an essential prerequisite for reliable and precise performance of downstream vision-based transportation tasks.
@@
As image dehazing is a typical ill-posed problem, early works exploit various prior assumptions to generate haze-free images [8, 9, 10, 11, 12]. The hand-crafted prior assumptions can regularize the solution space of linear inverse problems. However, predefined assumptions are only applied in specific scenarios, causing severe color shift and undesired artifacts in practice. Recently, numerous learning-based dehazing approaches have been proposed. Early efforts [13, 14, 15] develop convolutional neural networks (CNNs) to estimate the transmission maps or airlight, and generate sharp images via atmospheric scattering model (ASM) [16]. However, the cumulative error and unreliable prior assumptions often cause color distortions or haze residuals. Since then, considerable end-to-end dehazing approaches [17, 18, 19, 20, 21, 22] directly generate haze-free images from the corresponding hazy images. With the advantage of generative models, some dehazing methods adopt generative adversarial networks (GANs) [23, 24, 25, 26, 27] to regularize data distribution of dehazed images for haze removal. However, the training of GANs is unstable and susceptible to mode corruption.
Notably, natural image dehazing (NID) and remote sensing image dehazing (RSID) have several differences: First, RS images are taken from higher altitudes, so their surface information contains extensive small-scale texture details from land cover, vegetation, and urban buildings. Second, the atmospheric scattering model applied in NID does not work on RSID, due to the invariable scene depth in RS images [21]. Third, RS images are usually captured from multiple bands. That is, RSID needs to consider the spectral characteristics of recovered images, especially when dealing with multispectral or hyperspectral data. Fourth, compared with the datasets for NID, there are limited datasets for RSID due to high available cost [21], thus it is valuable to construct large RS hazy datasets based on RS imaging mechanism.
@@
Diffusion model is an emerging generative paradigm and successful in various low-level vision tasks [28, 29, 30]. With the superior modeling data distribution ability, diffusion models have achieved impressive performance in image generation. Moreover, the diffusion model is not subjected to unstable training and mode-collapse as GAN-based methods. Motivated by the strong generation ability of diffusion models, we leverage the conditional DDPM [31] for RSID. However, there still exist several problems that limit the dehazing performance of DDPM. First, DDPM only uses simple noise-estimation constraint to optimize the network while ignoring image-level refinement. That simple constraining noise may introduce chaotic content and undesired impurity into restored images. Second, DDPM cannot utilize the potential structure and statistical information in the forward iterative process, limiting its better sampling capacity. During early iterations in the diffusion model, the forward iterative process preserves more structural and color information than reverse sampling process. It may provide additional beneficial information for finer sampling results, especially for dense haze removal. Note the early iterations mean that time step $t$ in DDPM is smaller. As shown in Fig. 1(d), by exploiting the beneficial information in the forward iterative process, DDPM can better recover image details and color information. Third, although the patch-based DDPM enables size-agnostic processing and sampling acceleration, it yields evident block artifacts in that the patch-based strategy overlooks the global semantics of all patches. As shown in Fig. 1(e)), introducing the global compensation learning module is beneficial to eliminate boundary artifacts and produce more visually pleasing images.
@@
To address these issues, we propose a novel DDPM-based network for RSID, termed RSHazeDiff. Similar to the conditional DDPM, RSHazeDiff also adopts the degraded images as conditions to prevent content diversity with the stochastic sampling process. In addition, we develop a phased training strategy (PTS) to remedy the unsatisfactory results caused by simple noise-estimation constraint in the training phase. Concretely, we first train the diffusion model with the noise-estimation constraint, and then further optimize it with the reconstruction constraint. The reconstruction constraint aims to constrain the sampling results with its corresponding ground-truth images. Moreover, to take the most of latent structure and statistical information in the forward iterative process, we design a Fourier-aware iterative refinement (FIR) module. In the early sampling step (i.e., time step $t$ is small), FIR excavates the style/semantic information of the amplitude/phase components in Fourier domain for finer sampling results. Referring to the patch-based image processing in [32], we leverage the patch-based image processing strategy for unlimited-size image dehazing. However, separately dealing with each patch ignores the global dependency features of all patches, causing evident artifacts in the final result. In view of this, we utilize a global compensated learning module (GCL) to capture global dependency features. Comprehensive experiments verify the superiority of RSHazeDiff on both synthetic and real-world RS hazy images, especially in dense hazy scenes.
@@
  • The global compensated learning module is introduced to eliminate evident block artifacts from the patch-based image processing scheme. By capturing the global dependency features in RS images, GCL assures the semantic consistency of all patches.
@@ II Related Work
This section briefly reviews RSID algorithms and the diffusion model for image restoration, followed by the Fourier transform in deep learning.
@@ II-A Remote Sensing Image Dehazing
Prior-based methods. These methods [10, 11, 12] employ the hand-crafted priors to restore the dehazed images. For instance, Xu et al. [10] design an iterative dehazing method named IDeRS, which introduces the concept of “virtual depth” for evaluating Earth surface coverings. Shen et al. [11] propose a spatial-spectral adaptive method for RSID, which corrects the ambient light and transmission estimation from the spatial and spectral dimensions. Han et al. [12] can effectively remove the haze of RS hazy images by exploiting multiscale guided filtering to decompose the input images into base layers and detail layers for final output. Although these prior-based methods achieve impressive results, they rely on prior assumptions in specific scenes, making the dehazing results easily violated when the assumptions are not satisfied.
Learning-based methods. With the success of deep learning, researchers have focused on learning-based RSID algorithms [20, 27, 21, 33, 34, 35, 36]. Similar to natural image dehazing, some RSID methods explore deep neural networks to estimate haze parameters in ASM. For example, Chi et al. [34] develop a trinity model called Trinity-Net to incorporate prior knowledge into the gradient-guided Swin Transformer [37] for haze parameters estimation. The other works directly convert hazy images to haze-free images in an end-to-end manner. Among them, Li and Chen [20] design a coarse-to-fine dehazing network called FCTF-Net, in which the first stage exploits an encoder–decoder architecture to obtain coarse dehazing results and then refines them in second stage. Zhang and Wang [21] propose a dynamic collaborative inference learning framework (DCIL) to handle haze removal of dense RS hazy images. Zheng et al. [27] exploit an enhanced attention-guide GAN-based network named Dehaze-AGGAN to restore haze-free remote sensing images in an unsupervised manner. However, few works focus on applying the diffusion model to image dehazing task, whereas the advanced generation ability of diffusion model makes it possible to better recover haze-free images.
@@ II-B Diffusion Models for Image Restoration
Diffusion-based generative models [38, 31] progressively corrupt images by adding noise, and iteratively denoise from the noise distribution to sharp images. With the superior generative capability, diffusion-based models have achieved remarkable performance in image restoration, such as image super-resolution [28], inpainting [29], and deblurring [30]. Saharia et al. [28] adopt conditional diffusion model for image super-resolution and achieve more competitive results than GAN-based methods. With similar condition strategy, Saharia et al. [29] propose a unified framework to handle image-to-image translation tasks by incorporating distorted images as condition input. Rombach et al. [39] apply diffusion models in latent space to improve sampling efficiency. Considering that current diffusion model architectures mainly focus on fixed-sized image restoration, Ozan et al. [32] propose a patch-based diffusion model for unlimited-size image restoration. However, it fails to capture global information when operating locally on fixed-sized patches, which yields apparent block artifacts. To address this issue, we design a global compensated learning module to capture long-range context in frequency domain.
@@ II-C Fourier Transform in Deep Learning
Recently, the frequency information extracted through Fourier transform has attracted increasing attention in deep learning. A line of works utilizes it to improve representation ability of neural networks [40, 41, 42]. Among them, Wang et al. [41] reveal the relationship between the generalization of CNNs and the high-frequency components of images. Besides, Zhou et al. [42] propose deep Fourier up-sampling for multi-scale modeling, which achieves impressive performance on multiple computer vision tasks in a plug-and-play manner. The other line of work applies the Fourier transform to many computer vision tasks [43, 44, 45, 46, 47, 48]. Taking image restoration tasks as an example, Fuoli et al. [45] adopt Fourier domain losses to better recover the high-frequency content in image super-resolution. Yu et al. [46] explore the haze degradation property from a frequency domain perspective to remove haze. Motivated by the convolution theorem that reveals an equivalent property between spatial domain and frequency domain, Kong et al. [47] propose an effective frequency domain-based Transformer for image deblurring, with less computational complexity. Inspired by some observations of low-light noisy images in the frequency domain, Li et al. [48] design a cascaded network with Fourier Transform for low-light image enhancement. However, no works attempt to exploit the potential of the Fourier spectral components in diffusion models. To our knowledge, we are the first to take the Fourier spectral components as important prior knowledge during iterative sampling steps of diffusion model, better preserving image details and colors.
@@ IV-A Overview
Fig. 2 illustrates the overall architecture of RSHazeDiff, which consists of three main components: a local reverse denoising process, a global compensated learning module, and a feature fusion module. The local reverse denoising process leverages the Fourier-aware condition diffusion model to perform local patch-level restoration, and an input hazy image is randomly cropped with several patches as its conditions. The detailed training and sampling process of Fourier-aware conditional diffusion model is exhibited in Fig. 3, and its pseudocodes are depicted in Alg. 1 and Alg. 2. Notably, during inference we crop the whole image into overlapping patches as the input of Fourier-aware conditional diffusion model by using the grid-like arranged parsing scheme [32]. In turn, we merge all patches with the dictionary of their locations and adopt the mean estimated noise for overlapping pixels [32] (lines 8,9,11 in Alg. 2). GCL module adopts the Fast Fourier Block (FFB) as the basic unit of U-net architecture to capture global feature representations. The structure of FFB is shown Fig. 4. Finally, the feature fusion module fuses the global-local feature representations to produce the final dehazing results.
@@ IV-B Fourier-aware Iterative Refinement
DDPM-based image restoration approaches tend to tackle the forward diffusion process and reverse denoising process in sequence. The reverse process progressively sample with variable $\epsilon_{0}\left(x_{t},\bar{x},t\right)$ predicted in the forward process. Apart from that, the two processes are independent to a certain extent. Nevertheless, early iterations in the forward process can provide beneficial guided features to the sampling process. As depicted in [50], the amplitude and phase components in Fourier space correspond to low-level statistics and high-level semantic information of input images. This characteristic can be extended in image dehazing task. The amplitude spectrum contains more hazy degradation properties (e.g., illumination and contrast), while the phase spectrum preserves more texture structure of images. In early sampling steps (time step $t$ is small), dense hazy images suffer from severe structure and color information loss, whereas the forward iterative process preserves more structural and color information. Inspired by that, we design a Fourier-aware iterative refinement module, which excavates the structure and color guided representation from the forward process in Fourier domain to refine transitions of sampling process. The well-designed FIR module can help diffusion model to generate more natural and realistic images.
The detailed structure of FIR is illustrated in Fig. 3. Note that the $x_{0}^{(i)}$ represents $i$-th patch cropped from the clear image $x_{0}$. we transition step from a patch $x_{t-1}^{(i)}$ to $x_{t}^{(i)}$ through Eq. 1, and then sample from $\hat{x}_{t}^{(i)}$ to $\hat{x}_{t-1}^{(i)}$ through Eq. 4. Then, images $x_{t-1}^{(i)}$ and $\hat{x}_{t-1}^{(i)}$ are decomposed to amplitude and phase components in Fourier domain by Fast Fourier Transform (FFT) [51]. Next, we use the phase spectrum of $x_{t-1}^{(i)}$ and integrate the partial amplitude spectrum of $x_{t-1}^{(i)}$ into amplitude component of $\hat{x}_{t-1}^{(i)}$. Finally, we apply the inverse Fast Fourier Transform (IFFT) to transform the final fused results $\hat{x}_{t-1}^{(i)}$ back to spatial domain image. Given a single channel image $x\in\mathcal{R}^{H\times W}$, its Fourier transform $\mathcal{F}$ as a complex component can be expressed as: